# Simple Path Structural Encoding for Graph Transformers

### v1

@@ 5.2 Real-World Benchmarks
Note that replacing walks by path count is done at no additional cost as the number of trainable parameters remains unchanged. Two-sided Student’s t-tests are conducted to assess the significance of the obtained results. Finally, we compare with the following GNN methods: GCN (Kipf and Welling, 2016), GIN (Xu et al., 2019), GAT (Velickovic et al., 2017), GatedGCN (Bresson and Laurent, 2017) and GIN-AK+ (Zhao et al., 2021), along with the graph transformers: SAN (Kreuzer et al., 2021), Graphormer (Ying et al., 2021), EGT (Hussain et al., 2022), GraphViT (He et al., 2023), Exphormer (Shirzad et al., 2023) and Drew (Gutteridge et al., 2023). Results are reported in Table 1.
@@ 5.4 Limitations
In the proposed algorithm, the path count between two nodes is updated by comparing the value provided by each new DAG with the total currently stored in memory. The latter is then replaced by the maximum of the two values. This approach is necessary since paths are not enumerated because of memory constraints. The obtained counts therefore constitute lower bounds on the exact number of paths (see line 16 of Algorithm 3 in Appendix D). Failure cases can however arise because individual paths are not distinguished. For instance, consider the input graph illustrated in Figure 6. The total number of paths of length 4 between nodes 0 and 1 is two (blue and green). The only directed graphs that would allow to discover these paths simultaneously are not acyclic: our algorithm can simply count them one at a time.
This limitation is especially true in high-density graphs: in fact our method does not yield any statistically significant improvement on the dense CLUSTER dataset.
@@ 6 Related Work
Graph Transformers. Graph transformers are a class of models designed to process graph-structured data by leveraging self-attention, which provides a graph-wide receptive field (Vaswani, 2017; Ying et al., 2021), thereby avoiding common limitations of message-passing GNNs (Topping et al., 2022). Despite their advantages, graph transformers have not yet become the dominant architecture in graph learning, primarily due to the lack of unified structural and positional encodings that generalize well across diverse graph tasks (Rampášek et al., 2022). This is evident from the continued success of hybrid approaches that integrate message-passing with global self-attention (Wu et al., 2021; Rampášek et al., 2022; He et al., 2023; Choi et al., 2024), highlighting ongoing opportunities for architectural improvements. Promising research directions include: more flexible spectral attention mechanisms (Bo et al., 2023), models that mitigate the quadratic memory footprint of self-attention (Shirzad et al., 2023), and hierarchical encodings that enhance structural representations (Luo et al., 2024b).
Expressivity of Path-Based MPNNs. The expressivity of message-passing neural networks (MPNNs) on paths has been studied in relation to the Weisfeiler-Leman (WL) isomorphism test (Michel et al., 2023; Graziani et al., 2023). Notably, Graziani et al. (2023) demonstrated that iteratively updating node features via message-passing along paths originating from these nodes is strictly more expressive than the 1-WL test. However, these works consider only short path lengths, limiting their applicability to more complex graph structures.
Counting Paths and Cycles in Graphs. The problem of counting paths and cycles in graphs has been extensively studied over the past decades (Johnson, 1975; Alon et al., 1997; Flum and Grohe, 2004). Perepechko and Voropaev (2009) derived an explicit formula for counting paths of length up to 6 between two nodes, while Giscard et al. (2019) proposed an algorithm for counting paths of any length based on enumerating connected induced subgraphs. However, these methods become impractical for the large graphs and long path lengths considered in our work, necessitating the use of efficient approximate counting methods.
@@ 7 Conclusion
This work introduced Simple Path Structural Encoding (SPSE), a novel structural encoding method for graph transformers that leverages path counts instead of random walk probabilities. By providing a more expressive and structurally informative edge encoding, SPSE improves performance across various graph learning benchmarks. Our theoretical and experimental study shows that SPSE mitigates the limitations of random walk-based encodings while maintaining computational feasibility through an efficient approximation algorithm. While promising, SPSE’s applicability to extremely large-scale graphs and its interaction with different transformer architectures would require further exploration. Future research directions may also include optimizing its computational efficiency and extending its applicability to broader domains such as knowledge graphs and large social networks.
@@ Appendix A Additional Path Count Statistics and Training Hyperparameters
Typical hyperparameter values for path counts with the respective computing duration are reported in Table 2. The hyperparameter search was conducted on a reduced pool of graphs and root nodes to optimize the trade-off between exhaustivity and duration. The reported figures should therefore not be viewed as optimal, but rather give an indication of how they affect the path count duration. The latter ranges from under one hour for ZINC to a bit more than three days for CIFAR10 and the large PCQM4Mv2 dataset which contains 3.7M graphs. Once computed, path counts are stored and used for model training at no additional cost. In Table 2 are also reported values of $\alpha$, $\beta$ and $n$ training hyperparameters (see equation 4).
Appendix B Proofs
B.1 Proof of Proposition 1
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an even-length cycle graph, i.e. $|\mathcal{V}|=2n$ for some $n\in\mathbb{N}^{*}$, and $\mathcal{G^{\prime}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ a path graph such that $|\mathcal{V}^{\prime}|=2n+1$. Then given any pair of nodes $(i,j)$ in $\mathcal{G}$, there exits a pair of nodes $(i^{\prime},j^{\prime})$ in $\mathcal{G^{\prime}}$ such that $(i,j)^{\mathcal{G}}=_{\textsc{rw}}(i^{\prime},j^{\prime})^{\mathcal{G^{\prime}}}$.
@@
The proof is done by induction by considering landing probabilities of walks of length $k$ and $k+1$ for any given $k$ in $\mathbb{N^{*}}$. Do do that, we use the following notations:
@@ B.3 Proof of Proposition 3
Appendix C Tree Decompositions
The formal proof of the equivalence of the RWSE encodings in circular and linear graphs is given in Appendix B.1. Here we give a graphical sketch of the proof for the two examples considered in Figure 1. The tree decompositions of the graphs in Figure 1 are presented in Figure 7 for walks starting in node 0 and a maximum depth of 5. As nodes only have one or two neighbors, colors are used to represent the walk probabilities toward child nodes. Edges $(0,1)$ have the same random walk probabilities in (A) and (B) (Figure 7, top) on one hand, and in (C) and (D) (Figure 7, bottom) on the other hand, despite belonging to very different graphs. It is straightforward to extend these results to any depth, hence verifying that $(0,1)^{A}=_{\textsc{rw}}(0,1)^{B}$, and $(0,1)^{C}=_{\textsc{rw}}(0,1)^{D}$.
@@ Appendix C Tree Decompositions
Appendix D Full Path Counting Algorithm
@@
Appendix E Synthetic Experiment

### v2

@@ 5.2 Real-World Benchmarks
Note that replacing walks by path count is done at no additional cost as the number of trainable parameters remains unchanged. Two-sided Student’s t-tests are conducted to assess the significance of the obtained results. Finally, we compare with the following GNN methods: GCN (Kipf & Welling, 2017), GIN (Xu et al., 2019), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2017) and GIN-AK+ (Zhao et al., 2022), along with the graph transformers: SAN (Kreuzer et al., 2021), Graphormer (Ying et al., 2021), EGT (Hussain et al., 2022), GraphViT (He et al., 2023), Exphormer (Shirzad et al., 2023), Drew (Gutteridge et al., 2023) and GPSE (Cantürk et al., 2024). Results are reported in Table 1.
@@ 5.4 Limitations
In the proposed algorithm, the path count between two nodes is updated by comparing the value provided by each new DAG with the total currently stored in memory. The latter is then replaced by the maximum of the two values. This approach is necessary since paths are not enumerated because of memory constraints. The obtained counts therefore constitute lower bounds on the exact number of paths (see line 16 of Algorithm 3 in Appendix E). Failure cases can however arise because individual paths are not distinguished. For instance, consider the input graph illustrated in Figure 6. The total number of paths of length 4 between nodes 0 and 1 is two (blue and green). The only directed graphs that would allow to discover these paths simultaneously are not acyclic: our algorithm can simply count them one at a time.
This limitation is especially true in high-density graphs, such as the CLUSTER dataset where our method does not yield any significant improvement. This suggests that there may exist cases where inaccurate path counts are detrimental to the overall performance, and where it might be preferable to trade longer path lengths for exact counts on shorter paths.
@@ 6 Related Work
Graph Transformers. Graph transformers are a class of models designed to process graph-structured data by leveraging self-attention, which provides a graph-wide receptive field (Vaswani, 2017; Ying et al., 2021), thereby avoiding common limitations of message-passing GNNs (Topping et al., 2022). Despite their advantages, graph transformers have not yet become the dominant architecture in graph learning, primarily due to the lack of unified structural and positional encodings that generalize well across diverse graph tasks (Rampášek et al., 2022). This is evident from the continued success of hybrid approaches that integrate message-passing with global self-attention (Wu et al., 2021; Rampášek et al., 2022; He et al., 2023; Choi et al., 2024), highlighting ongoing opportunities for architectural improvements. Promising research directions include: more flexible spectral attention mechanisms (Bo et al., 2023), models that mitigate the quadratic memory footprint of self-attention (Shirzad et al., 2023), and hierarchical encodings that enhance structural representations (Luo et al., 2024a).
Expressivity of Path-Based MPNNs. The expressivity of message-passing neural networks (MPNNs) on paths has been studied in relation to the WL isomorphism test (Michel et al., 2023; Graziani et al., 2024). Notably, Graziani et al. (2024) demonstrated that iteratively updating node features via message-passing along paths originating from these nodes is strictly more expressive than the 1-WL test. However, these works consider only short path lengths, limiting their applicability to more complex graph structures.
Counting Paths and Cycles in Graphs. The problem of counting paths and cycles in graphs has been extensively studied over the past decades (Johnson, 1975; Alon et al., 1997; Flum & Grohe, 2004). Perepechko & Voropaev (2009) derived an explicit formula for counting paths of length up to 6 between two nodes, while Giscard et al. (2019) proposed an algorithm for counting paths of any length based on enumerating connected induced subgraphs. However, these methods become impractical for the large graphs and long path lengths considered in our work, necessitating the use of efficient approximate counting methods.
@@ 7 Conclusion
This work introduced Simple Path Structural Encoding (SPSE), a novel structural encoding method for graph transformers that leverages path counts instead of random walk probabilities. By providing a more structurally informative edge encoding, SPSE improves performance across various graph learning benchmarks. Our theoretical and experimental study shows that SPSE mitigates the limitations of random walk-based encodings while maintaining computational feasibility through an efficient approximation algorithm. While promising, SPSE’s applicability to extremely large-scale graphs and its interaction with different transformer architectures would require further exploration. In particular, SPSE bears particularly promising synergies with hierarchical methods, which would allow efficient path mining at higher hierarchy levels while capturing valuable long-range structural information. Future research directions may also include optimizing its computational efficiency and extending its applicability to broader domains such as knowledge graphs and large social networks.
Acknowledgments
This work was funded in part by a grant from Fondazione Caritro, under project GenIE. AL and AP acknowledge the support of the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU.
@@ Appendix A Additional Path Count Statistics and Training Hyperparameters
Typical hyperparameter values for path counts with the respective computing duration are reported in Table 2. The hyperparameter search was conducted on a reduced pool of graphs and root nodes to optimize the trade-off between exhaustivity and duration. The reported figures should therefore not be viewed as optimal, but rather give an indication of how they affect the path count duration. The latter ranges from under one hour for ZINC to a bit more than three days for CIFAR10 and the large PCQM4Mv2 dataset which contains 3.7M graphs. Once computed, path counts are stored and used for model training at no additional cost.
Path count encoding parameters $\alpha$, $\beta$ and $n$. In the absence of clear heuristics regarding the effects of parameters $\alpha$, $\beta$ and $n$ of Equation 4, we report in Table 2 the values that led to the best results for each dataset. Among those, $n$ primarily controls the compression of the dynamic range of path counts, while both $\alpha$ and $n$ contribute to the adjustment of the output range of function $f$ in Equation 4. Unsurprisingly, denser datasets call for a higher $n$, a lower $\alpha$, or a combination of both.
@@
Appendix B Expressivity of SPSE encoding
We provide here a discussion about the expressivity of the proposed SPSE encoding in light of findings from previous works. A reasoning similar to the one followed for Path-WL (Graziani et al., 2024) can be used to show that an iterative node coloring algorithm based on global attention with SPSE is more expressive than 1-WL. We note also that the results regarding the expressivity of GRIT remain valid when SPSE replaces RWSE (called RRWP therein) as the chosen structural encoding method, except for the case of Proposition 3.1 (b) which requires access to transition probabilities.
However, no previous study allows the comparison of SPSE and RWSE expressivity. Contrary to WL- / Path-trees used in Graziani et al. (2024) and Michel et al. (2023), SPSE and RWSE aggregate information over simple paths and walks without enumerating them, which prevents one from using the strategy of the proof of Theorem 3.3 of Michel et al. (2023) to conclude. We leave the task of proving whether both encodings can be compared for future work.
Appendix C Proofs
C.1 Proof of Proposition 1
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an even-length cycle graph, i.e. $|\mathcal{V}|=2n$ for some $n\in\mathbb{N}^{*}$, and $\mathcal{G^{\prime}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ a path graph such that $|\mathcal{V}^{\prime}|=2n+1$. Then given any pair of nodes $(i,j)$ in $\mathcal{G}$, there exists a pair of nodes $(i^{\prime},j^{\prime})$ in $\mathcal{G^{\prime}}$ such that $(i,j)^{\mathcal{G}}=_{\textsc{rw}}(i^{\prime},j^{\prime})^{\mathcal{G^{\prime}}}$.
@@
The proof is done by induction by considering landing probabilities of walks of length $k$ and $k+1$ for any given $k$ in $\mathbb{N^{*}}$. To do that, we use the following notations:
@@ C.3 Proof of Proposition 3
Appendix D Tree Decompositions
The formal proof of the equivalence of the RWSE encodings in circular and linear graphs is given in Appendix C.1. Here we give a graphical sketch of the proof for the two examples considered in Figure 1. The tree decompositions of the graphs in Figure 1 are presented in Figure 7 for walks starting in node 0 and a maximum depth of 5. As nodes only have one or two neighbors, colors are used to represent the walk probabilities toward child nodes. Edges $(0,1)$ have the same random walk probabilities in (A) and (B) (Figure 7, top) on one hand, and in (C) and (D) (Figure 7, bottom) on the other hand, despite belonging to very different graphs. It is straightforward to extend these results to any depth, hence verifying that $(0,1)^{A}=_{\textsc{rw}}(0,1)^{B}$, and $(0,1)^{C}=_{\textsc{rw}}(0,1)^{D}$.
@@ Appendix D Tree Decompositions
Appendix E Full Path Counting Algorithm
@@
Appendix F Synthetic Experiment